radomes.openskiesproject.org
00Jzlyne00 00Jzlyne00
  • 26-01-2015
  • Physics
Answered

Support that P(e)=0.76,P(F)=0.45,and P(E and F)=0.32. What is P (E or F)?

Verified Answer :

AL2006
AL2006 AL2006
  • 26-01-2015


The probability of (E and F) = P(e) times P(f) = 0.76 x 0.45 = 0.342 .

The probability of (e OR f) =

                                   1 - P(not e AND not f) =

                                   1 - P(not e) x P(not f) =

                                   1 - (1 - 0.76) x (1 - 0.32) =

                                   1 - ( 0.24 )   x   ( 0.68 )  =  1 - 0.1632  =  0.8368.


Answer Link

Related Questions

compare the reign of Louis XIV with that of Peter the great which absolute ruler had more lasting impact on his country explain why
what was the basic message jesus preached how and why did that message spread
what does 621k 2 7k 5 equal
H4X pKa1 15 H3X 2 pKa2 36 H2X 3 pKa3 78 HX 4 pKa4 109 X 5 pKa5 115 867 545 1345
How are spores structurally different from seeds
What does feminism mean
Douglass writes about the songs the slaves sang How does he react to this topic a He sighs because he misses hearing the songs b He is moved because the songs m
Simplify each expression: 14. $\frac{(-7)^{5-7}}{(-7) \div(-7)^4}$ 15. $\frac{8^2}{8^4}$ 19. $\frac{4^3 \cdot 4^0}{4^2 \cdot 4^3}$ 20. $\frac{9^{12}
The Igbo and Ghana were two societies that had a similar organization based on religion TRUE or FALSE
how would you describe enrique salaverria sculpture of a woman with a building raising up from her head